Let r R Q Pr 4 () p = be the charge density distribution for a solid sphere of radius R and total charge Q. For a point ‘p’ inside the sphere at distance r1 from the centre of the sphere, the magnitude of electric field is –

Grade:11

2 Answers

Arun
3 years ago
 
Dear Jayant
 
Please check and repost the question. You can also attach an image. I will be happy to help you.
Vikas TU
3 years ago
If the sphere is a conducting sphere, then the electric field inside will be Zero.
If the sphere is non conducting, then consider a gaussian surface to find the electric field at a point inside.
So, according to Gauss's Law,
E{4πr^2}=1/ε∫ρdv
⇒E=14πε0^2{∫Q/rπR^4[4πr^2dr]}
⇒E=4Q/4πε0R^4r^2{∫r^3dr}
⇒E=Q/πε0R^4r^2r^4/4
E=Qr^2/4πε0R^4